### v3

@@ Abstract
In this paper, we consider a class of non-convex and non-smooth sparse optimization problems, which encompass most existing nonconvex sparsity-inducing terms. We show the second-order optimality conditions only depend on the nonzeros of the stationary points. We propose two damped iterative reweighted algorithms including the iteratively reweighted $\ell_{1}$ algorithm (DIRL1) and the iteratively reweighted $\ell_{2}$ (DIRL2) algorithm, to solve these problems. For DIRL1, we show the reweighted $\ell_{1}$ subproblem has support identification property so that DIRL1 locally reverts to a gradient descent algorithm around a stationary point. For DIRL2, we show the solution map of the reweighted $\ell_{2}$ subproblem is differentiable and Lipschitz continuous everywhere. Therefore, the map of DIRL1 and DIRL2 and their inverse are Lipschitz continuous, and the strict saddle points are their unstable fixed points. By applying the stable manifold theorem, these algorithms are shown to converge only to local minimizers with randomly initialization when the strictly saddle point property is assumed.
@@ 1 Introduction
where $f:\mathbb{R}^{n}\rightarrow\mathbb{R}$ and the penalty/regularization term $\psi(x)$ takes the form $\psi(x):=\sum\limits_{i=1}^{n}s(x_{i})$ with $s(x_{i})=r(|x_{i}|)$. The data fitting function $f$ is twice continuously differentiable and possibly nonconvex, and $r$ satisfies the following properties.
@@
$r$ is continuous and concave on $[0,+\infty)$ with $r(0)=0$, C2-smooth on $(0,+\infty)$, $r^{\prime}(|0^{+}|)>0$ and $r^{\prime}(t)\geq 0$ for all $t>0$.
This assumption indicates $r(t)>0$ for all $t>0$, but $r$ could be nonconvex or not Lipschitz around 0.
The regularization term $\psi$ in (P) encourages sparse solutions, offering a favorable balance between model complexity, interpretability, and predictive performance. This makes it well-suited for a wide range of tasks including machine learning, statistics, signal processing, and image analysis [1, 2, 3, 4]. Many widely used nonconvex regularization such as the smoothly clipped absolute deviation penalty (SCAD) [5] or the minimax concave penalty (MCP) [6], the $\ell_{p}$ quasi-norm penalty [7] satisfy Assumption 1 and can offer advantages including improved sparsity and better model interpretability. We provide the explicit form of several commonly used nonconvex regularization in Table 1. It is worth noting that these regularizers can be categorized into two classes: (i) Lipschitz continuous regularizer with bounded subdifferential around 0 ($r^{\prime}(|0^{+}|)<\infty$); (ii) non-Lipschitz continuous regularizer with unbounded subdifferential around 0 ($r^{\prime}(|0^{+}|)=\infty$) such as $\ell_{p}$ quasi-norm (LPN).
@@
In smooth optimization, many methods have been developed and analyzed to escape from saddle points, including first-order methods [8, 9, 10], momentum-based methods [11], second-order optimization algorithms [12], the algorithmic frameworks for constrained nonlinear problems [13], stochastic algorithms [14, 15], and techniques explicitly designed to identify and circumvent saddle points during optimization, such as the random perturbation technique [16, 17, 18, 19]. In smooth settings, it is widely assumed that the strict saddle property holds, meaning all saddle points possess at least one strictly negative eigenvalue. This property intuitively implies that at any saddle point, there exists a direction in which the objective function strictly decreases and it has been shown to be generally satisfied for smooth problems [20, 21].
@@
In nonsmooth optimization, avoiding saddle points poses a greater challenge. At stationary points, traditional tools such as the (sub)differential may not reliably indicate descent directions or facilitate the avoidance of saddle points. Developing nonsmooth optimization algorithms capable of effectively converging to desirable solutions and circumventing saddle points is inherently more complex than in smooth optimization. An example of this complexity is demonstrated in the analysis of the Douglas splitting algorithm for nonsmooth problems, as discussed in [22]. Recent research [23] has made strides in defining saddle points in nonsmooth cases as those occurring on the (smooth) active manifold by assuming $0\in\text{rint}\partial F(x)$ is satisfied at the stationary point. Additionally, several proximal methods, including proximal point method, proximal gradient method and proximal linear method [23], as well as a more efficient algorithm incorporating perturbation [24], have been proposed to converge to local minima in nonsmooth optimization settings. In these studies, the nonsmooth objective function $r(x)$ is assumed to be $\rho$-weakly convex meaning $r(\cdot)+\frac{\rho}{2}\|\cdot\|^{2}$ is convex. They assume an active-manifold around the limit point and then define the strict saddle points as those of the $f$ restricted to the manifold. The weakly $\rho$-convexity for $r$ plays a crucial role in constructing C2 differentiable Moreau envelope with the same stationary points and strict saddle points in proximal methods.
@@
As for the regularized problem (P), several first-order methods have been proposed. [25] introduces a proximal gradient method tailored for solving the $\ell_{p}$ regularized problem with $10<p<$ 1. This method is subsequently extended to address group $\ell_{p,q}$ norm regularized problems in [26]. Furthermore, the Difference of Convex (DC) algorithm is enlisted [27] to handle nonconvex regularization problems, given that many existing nonconvex regularizers $r(x)$ can be reformulated as DC functions. It is important to note, however, that the DC algorithm generally assumes $r^{\prime}(|0^{+}|)<\infty$ to accommodate the DC decomposition, which does not hold true for the $\ell_{p}$ regularized problem. Another avenue of research involves iteratively reweighted algorithms, which approximate the nonconvex regularizers by iteratively reweighted $\ell_{1}$ or $\ell_{2}$ regularizers. Initially applied to solve the unconstrained $\ell_{p}$ regularized problem in [28] without convergence analysis, this approach was later employed to address general constrained nonconvex regularization problems [29]. Subsequent work in [30] established its global convergence and demonstrated its locally stable support property [31]. Furthermore, convergence rate analysis under the Kurdyka-Łojasiewicz property [32] and its extrapolated version was proposed in [33]. While these first-order methods primarily focus on identifying first-order stationary points, none of them guarantee the avoidance of saddle points.
@@ 1.1 The main results
In this paper, our main purpose is to develop and analyze first-order methods that can provably avoid strict saddle points of the nonconvex regularization problem (P) without the assumption of weakly convexity. In particular, we focus on iteratively reweighted methods which approximate the regularization term by simpler weighted $\ell_{1}$ or $\ell_{2}$ norms [34, 35, 36, 30, 37, 29, 31, 33]. In contrast to the proximal gradient algorithm, iteratively reweighted methods solve a sequence of weighted $\ell_{1}$ or $\ell_{2}$ regularized subproblems and are widely applicable to different non-convex regularization problems.
The iterative reweighted $\ell_{1}$ method (${\rm IRL}_{1}$) solves a sequence of weighted $\ell_{1}$ subproblems, which has an explicit solution characterized by the soft-threshold operator. This method offers a favorable local property [31], known as the model identification, under the following additional assumption.
@@
This kind of assumption is also widely made to guarantee that the existence of an active manifold around a stationary point as assumed by [23]. At the tail end of the iteration, the sign of the iterates remain unchanged while the nonzero components are kept uniformly bounded away from 0. In other words, the algorithm identifies a model subspace defined by an index set $\mathcal{J}$ as follows after finite number of iterations.
@@
(Model Subspace) The model subspace defined by $\mathcal{J}\subset\{1,\ldots,n\}$ is
$\mathcal{M}_{\mathcal{J}}:=\{u\in\mathbb{R}^{n}:\forall j\in\mathcal{J},u_{j}=0\}.$
This means every limit point of the iterates is also contained in $\mathcal{M}_{\mathcal{J}}$. This property motivates us to treat the mapping of ${\rm IRL}_{1}$ as smooth (locally around a limit point) on $\mathcal{M}_{\mathcal{J}}$, even though the underlying optimization problem is nonsmooth in $\mathbb{R}^{n}$.
The iterative reweighted $\ell_{2}$ method (${\rm IRL}_{2}$) is also widely applied to these non-convex regularization problems, which solves a sequence of weighted $\ell_{2}$ subproblems and the solution is trivially given. Compared with ${\rm IRL}_{1}$, ${\rm IRL}_{2}$ does not enjoy the model subspace identification/active-manifold identity property. In fact, it can be often witnessed that this algorithm can generate iterates with elements asymptotically approaching zero. To show the different behaviors of the two algorithms, we provide the following example.
@@
Let $g(x;x^{k})=\nabla f(x^{k})^{T}(x-x^{k})+\tfrac{1}{2}\|x-x^{k}\|^{2}$. The iterations of iteratively reweighted algorithms are as follows [29]. We use the ${\rm IRL}_{1}$ algorithm with iteration
@@
The behavior of ${\rm IRL}_{2}$ brings another challenge of analyzing the avoidance of strict saddle point. Current analysis for nonsmooth optimization algorithms ubiquitously depends on the active-manifold identification property, which ensures the algorithm reverts to a smooth algorithm in the active manifold in a local neighbourhood of the stationary point. However, there is no answer how a nonsmooth first-order algorithm like ${\rm IRL}_{2}$ would behave in the local region of a strict saddle point.
@@
    We first show the local minimal and second-order optimality of $f$ are equivalent to those of $f$ restricted on $\mathcal{M}_{\mathcal{J}}$, and define the strict saddle points as those on $\mathcal{M}_{\mathcal{J}}$.
@@
    We develop damped IRL1 algorithms individually for the two types of regularizers, $r^{\prime}(|0^{+}|)<\infty$ and $r^{\prime}(|0^{+}|)=\infty$, meaning the next iterate is the combination of a traditional IRL1 iterate and its previous iterate. The only difference between these two cases is that a relaxation $\epsilon>0$ is added to the regularizers for the latter case to prevent $r^{\prime}(\cdot)$ from becoming $\infty$ at each iteration. At the tail end of the iterations, we show this algorithm reverts to a damped gradient descent algorithm of the components in $\mathcal{J}$ on the subspace $\mathcal{M}_{\mathcal{J}}$ and decrease the rest components to 0 at a linear speed. We then apply the center stable manifold theorem and prove that the algorithm can avoid the convergence to strict saddle points.
@@
    We develop damped IRL2 algorithms for the nonconvex regularized problems and provide the global convergence. Under an additional assumption which is satisfied by most common nonconvex regularizers, we show that the algorithm is $C^{1}$ smooth and nondegenerate everywhere. We also proved the equivalence between the unstable fixed point of the iteration map and the strict saddle points of the regularization problem. Therefore, the avoidance of strict saddle points are derived. This is the first algorithm shown to have this property without using the active-manifold identification property nor the stochastic technique in nonsmooth optimization algorithms.
@@ 1.3 Notation
${\rm sign}(t)=\left\{\begin{aligned} &1,\ &t>0,\\ &-1,\ &t<0,\\ &0,\ &t=0.\end{aligned}\right.$ (1)
$DT(x)$ denote the Jacobian matrix of a map $T:\mathbb{R}^{n}\rightarrow\mathbb{R}^{n}$ at $x$. Given $x\in\mathbb{R}^{n}$ and $\mathcal{I}\subset\{1,\ldots,n\}$, denote $x_{\mathcal{I}^{*}}\subset\mathbb{R}^{\mathcal{I}^{*}}$ or $[x_{i}]_{i\in\mathcal{I}}$ as the subvector of the elements in $\mathcal{I}$; also, given $f:\mathbb{R}^{n}\rightarrow\mathbb{R}$ smooth with respect to $x_{i},i\in\mathcal{I}$, denote $\nabla_{\mathcal{I}}f(x)$ and $\nabla^{2}_{\mathcal{II}}f(x)$ as the gradient and the Hessian of $f$ at $x$, respectively, by fixing the values of $x_{i},i\in\{1,\ldots,n\}\setminus\mathcal{I}$ as constants. If $f$ is smooth, $\nabla^{2}_{\mathcal{II}}f(x)=[\nabla^{2}f(x)]_{\mathcal{II}}$ and also let $\nabla^{2}_{\mathcal{IJ}}=[\nabla^{2}f(x)]_{\mathcal{IJ}}$ for $J\subset\{1,\ldots,n\}$. Denote $f_{|\mathcal{I}}$ as the restriction of $f$ on $\mathcal{M}_{\mathcal{I}}$, i.e., by fixing $x_{i}\equiv 0,i\in\{1,\ldots,n\}\setminus\mathcal{I}$, $f_{|\mathcal{I}}(x_{\mathcal{I}}):=f([x_{\mathcal{I}^{*}},0_{\mathcal{J}^{*}}]).$ Let $I$ and $\bm{0}$ be the identity matrix and zero matrix of appropriate sizes, respectively. Given a symmetric matrix $H\in\mathbb{R}^{n\times n}$, let $\lambda_{\min}(H)$ be the smallest eigenvalue of $H$. Let $x\in\mathbb{R}^{n}$ and $w\in\mathbb{R}_{+}^{n}$, the soft-thresholding operator $\mathbb{S}_{w}:\mathbb{R}^{n}\to\mathbb{R}^{n}$ with the vector of thresholds $w$ is defined $[\mathbb{S}_{w}(x)]_{i}=\text{sign}(x_{i})\max(|x_{i}|-w_{i},0)$. In particular, if $w_{i}=\infty$, it follows that $[\mathbb{S}_{w}(x)]_{i}=0$.

### v4

@@ Abstract
In this paper, we consider a class of nonconvex and nonsmooth sparse optimization problems, which encompass most existing nonconvex sparsity-inducing terms. We show the second-order optimality conditions only depend on the nonzeros of the stationary points. We propose two damped iterative reweighted algorithms, including the iteratively reweighted $\ell_{1}$ algorithm (DIRL1) and the iteratively reweighted $\ell_{2}$ (DIRL2) algorithm, to solve these problems. For DIRL1, we show the reweighted $\ell_{1}$ subproblem has support identification property so that DIRL1 locally reverts to a gradient descent algorithm around a stationary point. For DIRL2, we show the solution map of the reweighted $\ell_{2}$ subproblem is differentiable and Lipschitz continuous everywhere. Therefore, the map of DIRL1 and DIRL2 and their inverse are Lipschitz continuous, and the strict saddle points are their unstable fixed points. By applying the stable manifold theorem, these algorithms are shown to converge only to local minimizers with random initialization when the strictly saddle point property is assumed.
@@ 1 Introduction
where $f:\mathbb{R}^{n}\rightarrow\mathbb{R}$ and the penalty/regularization term $\psi(x)$ takes the form $\psi(x):=\sum\limits_{i=1}^{n}s(x_{i})$ with $s(x_{i})=r(|x_{i}|)$. The data fitting function $f$ is differentiable twice continuously and possibly nonconvex, and $r$ satisfies the following properties.
@@
$r$ is continuous and concave on $[0,+\infty)$ with $r(0)=0$, C2 smooth on $(0,+\infty)$, $r^{\prime}(|0^{+}|)>0$ and $r^{\prime}(t)\geq 0$ for all $t>0$.
This assumption indicates $r(t)>0$ for all $t>0$, but $r$ could be nonconvex or not Lipschitz continuous around 0.
The regularization term $\psi$ in (P) encourages sparse solutions, offering a favorable balance between model complexity, interpretability, and predictive performance. This makes it well-suited for a wide range of tasks including machine learning, statistics, signal processing, and image analysis [1, 2, 3, 4]. Many widely used nonconvex regularization such as the EXP penalty [38] or the LOG penalty [39], the $\ell_{p}$ quasi-norm penalty [7] satisfy Assumption 1 and can offer advantages including improved sparsity and better model interpretability. We provide the explicit form of several commonly used nonconvex regularizations in Table 1. It is worth noting that these regularizers can be categorized into two classes: (i) Lipschitz continuous regularizer with bounded subdifferential around 0 ($r^{\prime}(|0^{+}|)<\infty$); (ii) nonLipschitz continuous regularizer with unbounded subdifferential around 0 ($r^{\prime}(|0^{+}|)=\infty$) such as the $\ell_{p}$ quasi-norm (LPN).
@@
In smooth optimization, many methods have been developed and analyzed to escape saddle points, including first-order methods [8, 9, 10], momentum-based methods [11], second-order optimization algorithms [12], the algorithmic frameworks for constrained nonlinear problems [13], stochastic algorithms [14, 15], and techniques explicitly designed to identify and circumvent saddle points during optimization, such as the random perturbation technique [16, 17, 18, 19]. In smooth settings, it is widely assumed that strict saddle property holds, which means that all saddle points have at least one strictly negative eigenvalue. This property intuitively implies that at any saddle point, there exists a direction in which the objective function strictly decreases, and it has been shown to be generally satisfied for smooth problems [20, 21].
@@
In nonsmooth optimization, avoiding saddle points poses a greater challenge. At stationary points, traditional tools such as the (sub)differential may not reliably indicate descent directions or facilitate avoidance of saddle points. Developing nonsmooth optimization algorithms capable of effectively converging to desirable solutions and circumventing saddle points is inherently more complex than in smooth optimization. An example of this complexity is demonstrated in the analysis of the Douglas splitting algorithm for nonsmooth problems, as discussed in [22]. Recent research [23] has made strides in defining saddle points in nonsmooth cases as those occurring on the (smooth) active manifold by assuming $0\in\text{rint}\partial F(x)$ is satisfied at the stationary point. Additionally, several proximal methods, including the proximal point method, the proximal gradient method, and the proximal linear method [23], as well as a gradient method incorporating random perturbation [24], have been proposed to converge to local minima in nonsmooth optimization settings. In these studies, the nonsmooth objective function $r(x)$ is assumed to be $\rho$ weakly convex, which means that $r(\cdot)+\frac{\rho}{2}\|\cdot\|^{2}$ is convex. They assumed an active manifold around the limit point and then defined the strict saddle points as those of the $f$ restricted to the manifold. The weakly $\rho$-convexity for $r$ plays a crucial role in constructing the C2 differentiable Moreau envelope with the same stationary points and strict saddle points in the proximal methods.
@@
As for the regularized problem (P), several first-order methods have been proposed. [25] introduces a proximal gradient method tailored for solving the $\ell_{p}$ regularized problem with $10<p<$ 1. This method is subsequently extended to address the $\ell_{p,q}$ group norm regularized problems of the group $\ell_{p,q}$ in [26]. Furthermore, the Difference of Convex (DC) algorithm is enlisted [27] to handle nonconvex regularization problems, given that many existing nonconvex regularizers $r(x)$ can be reformulated as DC functions. It is important to note, however, that the DC algorithm generally assumes $r^{\prime}(|0^{+}|)<\infty$ to accommodate the DC decomposition, which does not hold true for the $\ell_{p}$ regularized problem. Another avenue of research involves iteratively reweighted algorithms, which approximate the nonconvex regularizers by iteratively reweighted $\ell_{1}$ or $\ell_{2}$ regularizers. Initially applied to solve the unconstrained $\ell_{p}$ regularized problem in [28] without convergence analysis, this approach was later employed to address general constrained nonconvex regularization problems [29]. Subsequent work in [30] established its global convergence and demonstrated its locally stable support property [31]. Furthermore, convergence rate analysis under the Kurdyka-Łojasiewicz property [32] and its extrapolated version was proposed in [33]. While these first-order methods primarily focus on identifying first-order stationary points, none of them guarantee the avoidance of saddle points.
@@ 1.1 The main results
In this paper, our main purpose is to develop and analyze first-order methods that can provably avoid strict saddle points of the nonconvex regularization problem (P) without the assumption of weakly convexity. In particular, we focus on iteratively reweighted methods which approximate the regularization term by simpler weighted $\ell_{1}$ or $\ell_{2}$ norms [34, 35, 36, 30, 37, 29, 31, 33]. In contrast to the proximal gradient algorithm, iteratively reweighted methods solve a sequence of weighted $\ell_{1}$ or $\ell_{2}$ regularized subproblems and are widely applicable to various nonconvex regularization problems.
The iterative reweighted $\ell_{1}$ method (${\rm IRL}_{1}$) solves a sequence of weighted $\ell_{1}$ subproblems, which has an explicit solution characterized by the soft-threshold operator. This method offers a favorable local property [31], known as the model/support identification, under the following additional assumption.
@@
This kind of assumption is also widely made to guarantee that the existence of an active manifold around a stationary point as assumed by [23]. At the tail end of the iteration, the sign of the iterates remain unchanged while the nonzero components are kept uniformly bounded away from 0. In other words, the algorithm identifies a model subspace defined by an index set $\mathcal{I}$ as follows after a finite number of iterations.
@@
(Model Subspace) The model subspace defined by $\mathcal{I}\subset\{1,\ldots,n\}$ is
:=$\mathcal{M}_{\mathcal{I}}:=\{u\in\mathbb{R}^{n}:\forall j\in\mathcal{I}^{c},u_{j}=0\}.$
This means every limit point of the iterates is also contained in $\mathcal{M}_{\mathcal{I}}$. This property motivates us to treat the mapping of ${\rm IRL}_{1}$ as smooth (locally around a limit point) in $\mathcal{M}_{\mathcal{I}}$, although the underlying optimization problem is nonsmooth in $\mathbb{R}^{n}$.
The iterative reweighted $\ell_{2}$ method (${\rm IRL}_{2}$) is also widely applied to these nonconvex regularization problems, which solves a sequence of weighted $\ell_{2}$ subproblems and the solution is trivially given. Compared to ${\rm IRL}_{1}$, ${\rm IRL}_{2}$ does not enjoy the model subspace identification/active-manifold identity property. In fact, it can be often witnessed that this algorithm can generate iterates with elements asymptotically approaching zero. To show the different behaviors of the two algorithms, we provide the following example.
@@
Let $g(x;x^{k})=\nabla f(x^{k})^{T}(x-x^{k})+\tfrac{1}{2}\|x-x^{k}\|^{2}$. The iterations of the iteratively reweighted algorithms are as follows [29]. We use the ${\rm IRL}_{1}$ algorithm with iteration
@@
The behavior of ${\rm IRL}_{2}$ brings another challenge in analyzing the avoidance of strict saddle point. Current analysis for nonsmooth optimization algorithms ubiquitously depends on the active-manifold identification property, which ensures the algorithm reverts to a smooth algorithm in the active manifold in a local neighbourhood of the stationary point. However, there is no answer how a nonsmooth first-order algorithm like ${\rm IRL}_{2}$ would behave in the local region of a strict saddle point.
@@
    We first show the local minimal and second-order optimality of $f$ are equivalent to those of $f$ restricted on $\mathcal{M}_{\mathcal{I}}$, and define the strict saddle points as those on $\mathcal{M}_{\mathcal{I}}$.
@@
    We develop damped IRL1 algorithms individually for the two types of regularizer, $r^{\prime}(|0^{+}|)<\infty$ and $r^{\prime}(|0^{+}|)=\infty$, meaning that the next iterate is the combination of a traditional IRL1 iterate and its previous iterate. The only difference between these two cases is that a relaxation $\epsilon>0$ is added to the regularizers for the latter case to prevent $r^{\prime}(\cdot)$ from becoming $\infty$ at each iteration. At the tail end of the iterations, we show that this algorithm reverts to a damped gradient descent algorithm of the components in $\mathcal{I}$ on the subspace $\mathcal{M}_{\mathcal{I}}$ and decrease the rest components to 0 at a linear speed. We then apply the center stable manifold theorem and prove that the algorithm can avoid the convergence to strict saddle points.
@@
    We develop damped IRL2 algorithms for nonconvex regularized problems and provide global convergence. Under an additional assumption which is satisfied by most common nonconvex regularizers, we show that the algorithm is $C^{1}$ smooth and nondegenerate everywhere. We also proved the equivalence between the unstable fixed point of the iteration map and the strict saddle points of the regularization problem. Therefore, the avoidance of strict saddle points is derived. This is the first algorithm shown to have this property without using the active-manifold identification property or the stochastic technique in nonsmooth optimization algorithms.
@@ 1.3 Notation
We denote $\|\cdot\|_{p}$ as the quasi-norm $\ell_{p}$ with $p\in(0,1)$, that is, $\|x\|_{p}=(\sum_{i=1}^{n}|x_{i}|^{p})^{\frac{1}{p}}$. This does not define a proper norm because of the lack of subadditivity. For $x\in\mathbb{R}^{n}$, $x_{i}$ denotes the $i$-th element of $x$. The index set of nonzero and zero components of $x$ is defined as $\mathcal{I}(x)=\{i\mid x_{i}\neq 0\}$, and $\mathcal{J}(x)=\{i\mid x_{i}=0\}$, respectively. Moreover, $\mathcal{I}^{*}$ and $\mathcal{J}^{*}$ denotes $\mathcal{I}(x^{*})$ and $\mathcal{J}(x^{*})$ for short. ${\rm diag}(x)$ is the diagonal matrix by placing elements of vector $x$ on the main diagonal. ${\rm sign}(x):\mathbb{R}^{n}\rightarrow\mathbb{R}^{n}$ is the sign function defined by
$[{\rm sign}(x)]_{i}=\left\{\begin{aligned} &1,\ &x_{i}>0,\\ &-1,\ &x_{i}<0,\\ &0,\ &x_{i}=0.\end{aligned}\right.$ (1)
$DT(x)$ denote the Jacobian matrix of a map $T:\mathbb{R}^{n}\rightarrow\mathbb{R}^{n}$ at $x$. Given $x\in\mathbb{R}^{n}$ and $\mathcal{I}\subset\{1,\ldots,n\}$, denote $x_{\mathcal{I}^{*}}\subset\mathbb{R}^{\mathcal{I}^{*}}$ or $[x_{i}]_{i\in\mathcal{I}}$ as the subvector of the elements in $\mathcal{I}$; also, given $f:\mathbb{R}^{n}\rightarrow\mathbb{R}$ smooth with respect to $x_{i},i\in\mathcal{I}$, denote $\nabla_{\mathcal{I}}f(x)$ and $\nabla^{2}_{\mathcal{II}}f(x)$ as the gradient and the Hessian of $f$ at $x$, respectively, by fixing the values of $x_{i},i\in\{1,\ldots,n\}\setminus\mathcal{I}$ as constants. If $f$ is smooth, $\nabla^{2}_{\mathcal{II}}f(x)=[\nabla^{2}f(x)]_{\mathcal{II}}$ and also let $\nabla^{2}_{\mathcal{IJ}}f(x)=[\nabla^{2}f(x)]_{\mathcal{IJ}}$ for $J\subset\{1,\ldots,n\}$. Denote $f_{|\mathcal{I}}$ as the restriction of $f$ on $\mathcal{M}_{\mathcal{I}}$, i.e., by fixing $x_{i}\equiv 0,i\in\{1,\ldots,n\}\setminus\mathcal{I}$, $f_{|\mathcal{I}}(x_{\mathcal{I}}):=f([x_{\mathcal{I}^{*}},0_{\mathcal{J}^{*}}]).$ Let $I$ and $\bm{0}$ be the identity matrix and zero matrix of appropriate sizes, respectively. Given a symmetric matrix $H\in\mathbb{R}^{n\times n}$, let $\lambda_{\min}(H)$ be the smallest eigenvalue of $H$. Let $x\in\mathbb{R}^{n}$ and $w\in\mathbb{R}_{+}^{n}$, the soft-thresholding operator $\mathbb{S}_{w}:\mathbb{R}^{n}\to\mathbb{R}^{n}$ with the vector of thresholds $w$ is defined $[\mathbb{S}_{w}(x)]_{i}=\text{sign}(x_{i})\max(|x_{i}|-w_{i},0)$. In particular, if $w_{i}=\infty$, it follows that $[\mathbb{S}_{w}(x)]_{i}=0$.